OpenStudy (anonymous):

Think of a five-digit number and call it N. When N has a 7 placed in front of it, the number formed is 5 times as large as when N has a 7 placed after it. What number is N?

OpenStudy (anonymous):

Placing a 7 in front of a 5 digit number is the same as adding 700000 to it. Placing a 7 after it is the same as multiplying it by 10 and adding 7.

OpenStudy (nurali):

N + 700,000 = 5 * (N*10+7) Simple algebra: thus N = 14285

OpenStudy (anonymous):

thhanks :)

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

this is a much easier number theory question than the last one you posted =]